# CaloHadronic: a diffusion model for the generation of hadronic showers

### v1

@@ F.1 Classifier Scores
We further compare the showers generated by the model to the Geant4 simulation by training a fully connected high-level classifier to distinguish between model generated and Geant4 simulated showers. The 5 input shower observables are the three center of gravity variables, the number of hits and energy ratio (total visible energy divided by the incident energy). 150k Geant4 showers and 500k showers generated by CaloHadronic were used. A standard scaling of the inputs is applied. A $60\%$ , $20\%$, $20\%$ data split is applied for training set, validation and test set, respectively. The classifier is implemented as a fully connected neural network with three layers (containing 8, 8, 1 nodes respectively) with LeakyReLU [85] activation functions. It is trained with the Adam optimizer [76] for 100 epochs using a binary cross-entropy loss. The final model epoch is chosen based on the lowest validation loss.
@@ Appendix G Resolution and Linearity
Figure 22 shows that the resolution of CaloHadronic is slightly worse than Geant4, while the linearity plot shows good agreement between the two. Fitting a function of the form $a+b/\sqrt{E_{0}}+c/E_{0}$ to the incident energy ($E_{0}$) versus relative resolution ($\sigma_{90}/\mu_{90}$) distribution yields the following parameters for Geant4 (CaloHadronic): a: $0.063\pm 0.003$ ($0.069\pm 0.005$), b: $0.16\pm 0.04$ ($0.40\pm 0.06$), c: $0.16\pm 0.04$ ($-0.2\pm 0.1$). However, it should be noted that these values cannot be directly interpreted as resolution components, as they average over different materials and thicknesses.

### v2

@@ F.1 Classifier Scores
We further compare the showers generated by the model to the Geant4 simulation by training a fully connected high-level classifier to distinguish between model generated and Geant4 simulated showers. The 5 input shower observables are the three center of gravity variables, the number of hits and energy ratio (total visible energy divided by the incident energy). 150k Geant4 showers and 500k showers generated by CaloHadronic were used. A standard scaling of the inputs is applied. A $60\%$ , $\%, 2020\%$ data split is applied for training set, validation and test set, respectively. The classifier is implemented as a fully connected neural network with three layers (containing 8, 8, 1 nodes respectively) with LeakyReLU [57] activation functions. It is trained with the Adam optimizer [48] for 100 epochs using a binary cross-entropy loss. The final model epoch is chosen based on the lowest validation loss.
@@ Appendix G Resolution and Linearity
Figure 22 shows that the resolution of CaloHadronic is slightly worse than Geant4, while the linearity plot shows good agreement between the two. Fitting a function of the form $a+b/\sqrt{E_{0}}+c/E_{0}$ to the incident energy ($E_{0}$) versus relative resolution ($\sigma_{90}/\mu_{90}$) distribution yields the following parameters for Geant4 (CaloHadronic): a: $0.063\pm 0.003$ ($0.069\pm 0.005$), b: $0.16\pm 0.04$ ($0.40\pm 0.06$), c: $0.16\pm 0.04$ ($0.2\pm 0.1$). However, it should be noted that these values cannot be directly interpreted as resolution components, as they average over different materials and thicknesses.